# A new Angle on Benchmarking Noncovalent Interactions

### v1

@@ 1 Introduction
Theoretical predictions have become an invaluable tool in modern science. As accurate benchmarks in quantum chemistry become more and more common and numerous, some questions have arisen about the applicability of the underlying benchmark methods. For increasingly large systems, there are growing discrepancies1 between the interaction energies of large molecules for the main benchmark method of quantum chemistry, which is CCSD(T) (coupled cluster2 including all single and double substitutions and quasiperturbative triple excitations3) and the main benchmark method of solid state materials science, which is QMC (quantum Monte Carlo).4, 5, 6, 7.
@@
The polarizability will increase with the system size of aromatic molecules, and so will the strength of their van der Waals interactions with each other. This polarizability may thus be increasingly hard to calculate with progressively larger basis sets. For example, even small molecules and atoms with huge polarizabilities need many diffuse functions in order to accurately describe these quantities.8
@@
Larger intermolecular interactions are possibly overestimated by the use of perturbative triples in the (T) part, as second-order Møller-Plesset perturbation theory (MP2) strongly overestimates9 intermolecular interactions of aromatic molecules. Various papers have proffered this argument, showing that post-CCSD(T) contributions should become relevant at some point.10, 11, 12 However, one has to bear in mind that while MP2 underestimates interaction energies of stacks, MP3 overcorrects (which has inspired the MP2.5 method13). CCSD can be seen as a summation of MP singles, doubles, and disconnected products thereof to infinite order, and underestimates interaction energies as it neglects connected triples (T), which make their first appearance in fourth order. In contrast, however, other post-Hartree-Fock or post-DFT methods such as RPA (random phase approximation14) and DFT-SAPT (density functional theory-based symmetry-adapted perturbation theory15) are argued to yield better, if not correct intermolecular interaction energies of larger dimers.10, 11
@@
For intermolecular interactions with medium to large basis sets, the exact basis set limit is usually bracketed by calculations excluding and including counterpoise corrections,16, 17 with very few notable exceptions, like, for example, the formic acid dimer.18 Thus, an increasingly large gap between these two calculations (including and excluding a basis set correction) would indicate a volatility of a certain post-Hartree-Fock method. Basis set incompleteness errors can also often be estimated using localized orbital or other reduced-scaling approaches.
@@
While the use of CCSD(T) gradients is very often computationally prohibitive, several structures can be considered in order to get a more complete picture of the interaction. This is usually done by just varying the distances, e.g. extending the S66 set19 of intermolecular interactions to the S66x8 set,20 which considers each complex at eight different multiples of the intermonomer separation (at fixed intramonomer geometries), for a total of 528 data points.
@@
Post-Hartree-Fock methods such as coupled cluster theory2 have a clear hierarchy. For example, when static correlation effects are present and not captured well by CCSD(T), one may attempt to walk up the cluster expansion staircase to CCSDT and CCSDT(Q), which will yield some insight about whether ‘the gold standard’ is really that golden.21, 22, 23
@@ 2 Computational Details
The calculations have been performed with various programs, depending on the method employed. The density functional theory (DFT) 24, 25 geometry optimizations utilizing the B97M-V 26 functional as well as the DLPNO (Domain-Based Local Pair Natural Orbital)-CCSD(T) 27, 28 calculations have been performed with ORCA 29 , DFT-SAPT30, 31 with MOLPRO32, MP333 with QChem34 because of their RI implementation35, and all dRPA14, 36, LNO (Local Natural Orbitals)-CCSD(T)37, LNO-CCSDT(Q)38, and canonical CCSD(T)3 / CCSDT(Q)39 calculations with the MRCC40 program suite. All LNO-CCSD(T) calculations have been using their respective MP2 basis set corrections as implemented in the MRCC program. While some canonical post-CCSD(T) calculations were also carried out using MRCC, the most demanding ones were performed using a development version of the CFOUR program system41.
@@
For the L7 (i.e., seven large noncovalent complexes) set of systems,42 geometries from the original reference have been utilized as-is. All other structures were optimized with the B97M-V functional26 using the QZVPP43 basis set with full symmetry. For example, in the case of the benzene dimer, the D6h symmetric saddle point has been utilized. This implies that all structures are directly on top of each other, as only their interaction energy matters. rewrite this word salad They are not parallely displaced, such as the minimum structures of the benzene44 and coronene dimers45 — they are not minima at all. Despite this, the monomer geometries have either been fixed or relaxed, staying within their respective symmetry, for the ethylene stacks, whereas for the acene stacks, we have employed completely relaxed geometries. The rationale for this approach is to be able to use as much symmetry as possible, since only the slope of interaction energies of increasingly large dimers matters.
@@
Throughout this study, Dunning’s correlation consistent basis sets with and without diffuse functions, denoted cc-pVZ46 and aug-cc-pVZ47, have been put to use and will sometimes be abbreviated by DZ, TZ, QZ, and 5Z. The notation {T,Q}Z refers to basis set extrapolation from TZ and QZ basis sets according to the familiar extrapolation formula.48
@@
Unless specifically indicated otherwise, interaction energies have been counterpoise (CP) corrected by the standard Boys and Bernardi procedure.49
@@
In order to shed additional light on the performance of CCSD(T)3 vs. fully iterative CCSDT50 and CCSDT(Q),39 we carried out additional calculations using two post-CCSD(T) methods that still only have CPU time scaling with system size N: CCSD(T)Λ 51, 52 and CCSDT-253, 50, 54. Cremer and coworkers55 presented a detailed perturbation theoretical analysis of the differences between full CCSDT and various approximations to it. The CCSDT-3 method omits the coupling term from the full CCSDT amplitudes equations, while CCSDT-2 additionally omits all coupling terms from the said equations. Both reduce the CPU time scaling from to at the expense of an energy difference starting at fifth order in perturbation theory:
@@
where it should be noted55 that is the Hermitian conjugate of .
@@ 3.1 L7 Set and the discrepancies between two reference methods
The original starting point of this study was, in fact, the discrepancy between the two reference methods CCSD(T) and FN-DMC (fixed node diffusion Monte Carlo). For this purpose, it is useful to reiterate the differences between them. In addition, we have also performed DFT-SAPT calculations for all these systems. According to the theoretical background of references 10, 11 and point one above, DFT-SAPT may be even more accurate than CCSD(T) for the interaction energies of ever larger systems. In Table 1, we show the DFT-SAPT together with the literature values of various post-Hartree-Fock methods and FN-DMC for four molecules of the L7 set42 as well as the buckycatcher.
@@
| Benzene2 | 10.01 | 7.219 | 6.159 | 11.860 | 11.21 | 13.419 | 19.719, 19.759 |
| GCGC | 52.01 | 36.042 | 35.759 | 54.7 | 57.01 | 56.142 | 74.742, 73.6,61 74.959 |
| C3A | 62.41 | 34.142 | 47.859 | 62.7 | 69.11 | 74.742 | 113.042, 109.2,61 109.759 |
| C2C2PD | 75.61 | 27.742 | 58.259 | 81.5 | 86.11 | 95.442 | 159.942, 154.3,61 155.759 |
| C3GC | 101.31 | 61.842 | 80.859 | 104.6 | 120.01 | 127.242 | 188.842, 182.2,61 183.759 |
| C60CPPA | 130.51 | 93.859 | 157.8 | 174.61 | 354.959 |
@@
It is apparent from these numbers that MP2 often overestimates the interaction energies of large conjugated molecules, whereas MP3 underestimates them. MP2.5,62 which relies on compensation between the errors of MP2 and MP3, still somewhat overestimates the interaction energies when going to larger molecules.63 The values obtained by DFT-SAPT are pretty much exactly between CCSD(T) and fixed-node diffusion Monte-Carlo (FN-DNC) values.1 If we plot interaction energies vs. number of atoms, we obtain values ranging from 0.79 (MP3) to 0.94 (FN-DMC). The slope of FN-DMC is with 1.09 kJ/mol per atom much smaller than the slope of CCSD(T) with 1.46 kJ/mol per atom- with DFT-SAPT and also MP2.5 (1.28 and 1.34 kJ/mol per atom, respectively) in this case being somewhat smaller than CCSD(T).
@@ 3.2 Intermolecular Interaction Energies of Larger Molecules
The first question is: Can we estimate these numbers also from interaction energy slopes of other molecules? In the case that they do, they may offer a hint about the accuracy of different post-Hartree-Fock methods for the interaction energies of large molecules. The systems that may come in mind are the benzene, coronene, and circumcoronene dimers, as the coronene dimer was already present in the original L7 set. The parallel displaced coronene dimer is yielding a significant difference between FN-DMC with an interaction energy of 75.6 kJ/mol and CCSD(T) with an interaction energy of 86.1 kJ/mol, which is more than 10 kJ/mol. Given, however, that the circumcoronene dimer already has 144 atoms, we need other systems as touchstones for our hypothesis: that we can an interpolate the behavior for really large systems from the slopes of interaction energies of polyaromatic dimers.
To stay as close to the aromatic behavior that we see in the above sequence, we chose the acene dimer series, thus again beginning with the interaction energies of the benzene through hexacene dimers. Starting from heptacene dimer, however, the monomers will undergo cyclodimerization,64, 65, 66 which means that if we want to relax geometries like in the L7 paper (even with symmetry constraints) we will have to stop at hexacene dimer. The latter, with 82 atoms, is about the same size as coronene dimer.
@@ 3.3 Basis Set Effects
First, we try to evaluate the basis set effects of the different methods. If, for example, polarization would be a major issue, the dissociation energies along a chain dimer sequence would not be on a straight line, but rather be flat, as the increasing polarization would be less well described by a small number of basis functions. The number of basis functions, however, also increases linearly with system size, yielding almost perfectly straight lines for different acene series, as shown in Figure 1a-c for LNO-CCSD(T) using tight settings. As counterpoise corrections for LNO and similar methods in small molecules are problematic owing to screening out a substantial part of the ”ghost atoms” contribution [see, e.g., Ref.67], we had to exclude the first couple of cp-corrected interaction energies (for example, the benzene and anthracene dimer for the DZ basis set and the benzene, anthracene and naphthalene dimer for the TZ basis set) in each graph.
@@ 3.4 Local Correlation
Having established the basis set effects, we can estimate canonical CCSD(T) at the basis set limit. Furthermore, especially with the advent of using local correlation methods and extrapolating them,68, 69, 70, 71 we can evaluate their accuracy when going to larger molecules. For this purpose, we compare these methods, especially for a small basis set of double-zeta quality. Again, all of the investigated series exhibit nearly perfect linear regression with an value of 0.996 for methods.
@@ 3.5 Electronic Structure Methods
Here, we investigate several post-Hartree-Fock methods, such as MP2, MP2.5, MP3, the dispersion contribution to DFT-SAPT, as well as dRPA. From these methods, DFT-SAPT and dRPA in particular are deemed to be as more accurate than perturbation theory for large molecules.10, 11 For the acene series, we obtain the trend of MP3 CCSD DFT-SAPT dRPA CCSD(T) MP2.5 MP2. This is hardly surprising, as MP2 very much overestimates the interaction energies of large aromatic molecules, with MP3 underestimating them, as does CCSD. Of course, the question arises about the accuracy of CCSD(T), dRPA, and DFT-SAPT.
For the ethylene stack series, we obtain the trend CCSD MP3 DFT-SAPT dRPA CCSD(T) MP2.5 MP2. As observed in the acene series, MP2 overestimates dispersion contributions while both CCSD and MP3 underestimate them. dRPA yields results similar to CCSD(T), which is clearly observed for the cc-pVZ (X=D,T,Q) basis sets. (We note in passing72 that RPA is equivalent to a ring-diagram simplification of coupled cluster with all doubles, CCD.)
@@ 3.6 Effects of the Geometry
So far, we have used geometries for both acenes and ethylene stacks in which the distance between the two monomers were fully optimized within its (high) symmetry. In order to show that the linear behavior is not an artefact, it may be worth to investigate another geometry with a different intermonomer distance. Here, we chose the stacks at the rather large optimized distance of the ethylene dimer, as shorter distances (e.g. at the hexatetraene distance) would result in a positive interaction energy for the ethylene dimer. The ethylene stack series with relaxed geometry, as discussed in the basis set subsection, exhibits a slope of 8.70 0.27 kJ/subunit. In contrast, the same series with the fixed geometry at the ethylene dimer distance of 4.46 Å shows a much smaller slope of 1.26 0.24 kJ/subunit, where the uncertainty is about one-fifth of the actual value.
@@ 3.7 Best Estimates of Slopes
Usually, the best estimates in energies are obtained by going to ever larger basis sets and then just extrapolating the energies to the complete basis set (CBS) limit. In our case, however, we can either calculate ever larger systems along the sequences studied with a smaller basis set or increase the basis set size thus adding one dimension to the problem. Our results are summarized in Tables 2 and 3. The slopes obtained from just the first two acene dimers, namely, benzene2 and naphthalene2, barely differ compared to those of the whole series up to pentacene; the cp-corrected DZ CCSD(T) slope changes by 0.26 kJ/subunit, and the cp-uncorrected DZ CCSD(T) slope even less at 0.17 kJ/subunit. In general, including larger systems in the fits increases the slopes. From these values, including the cc-pV(T,Q)Z basis set extrapolation which is still feasible for canonical CCSD(T) for the benzene and naphthalene dimer, we can estimate the canonical CCSD(T)/cbs values for the benzene-pentacene series, which will be -23.14-0.26=-23.40 kJ/subunit for the cp-corrected CCSD(T) results and -23.09-0.17=-23.26 kJ/subunit for cp-uncorrected CCSD(T). Finally, this values can then be extended to the hexacene dimer with an aug-cc-pV(Q,5)Z basis set extrapolation.
@@
Of more interest are the post-CCSD(T) slopes. The CCSDT and CCSDT(Q) results for acene dimers (benzene and naphthalene dimers) included in Tables 2 and 3 were obtained using rank-reduced coupled cluster methods described in Refs. 73, 74, 75. In these calculations, the coupled cluster equations are solved within a certain excitation subspace instead of the full space as in the canonical methods. This allows for reducing the costs of the calculations, but introduces an additional parameter (size of the excitation subspace) as a parameter in the rank-reduced formalism. Therefore, it is important that the results are stable with respect to the adopted value of this parameter. In the Supporting Information, we provide a detailed analysis of the convergence of the rank-reduced results with respect to size of the excitation subspace for benzene dimer and naphthalene dimer, showing that the numerical values reported here are sufficiently well converged for the present purposes. The only additional approximation employed in the rank-reduced calculations is the Cholesky decomposition of two-electron integrals. We used the full pivoting variant of this decomposition with a VeryTight threshold () for the diagonal elements. The errors resulting from the decomposition with these settings impact the calculated interaction energies by no more than a few thousands of kJ/mol. The CCSDT method reduces the slopes (in absolute value) by a sizeable 0.75 (cp-corrected DZ) and 0.78 (standard DZ) kJ/subunit. However, CCSDT(Q) (again, in absolute value) reverts this by 0.4 in both instances, shrinking the CCSD(T) to CCSDT(Q) differences in the slopes to 0.35 (cp-corrected DZ) and 0.38 (uncorrected DZ) kJ/subunit.
@@
For the Hartree-Fock exchange part, this is rather simple, as all calculations were performed at the largest basis sets and the cp-corrected and cp-uncorrected slopes are 11.95 and 11.96 kJ/subunit, respectively, yielding exchange-correlation slopes of -10.5 0.2 kJ/mol per acene ring dimer. For the largest acene dimer hexacene, this would make an interaction energy of approximately 63.0 kJ/mol, considering our best estimates. This value, however, is neglecting the intercepts of 3.3 for Hartree-Fock exchange and 1.9 for correlation, arriving at 57.8 1.4 kJ/mol. CCSD(T), for comparison, would yield slightly larger slopes of -11.1 and -10.9, with an interaction energy of 66.0 for the hexacene dimer. Including the intercepts (3.3 and 2.0), we arrive at a final CCSD(T) energy of 60.7 kJ/mol 1.5 kJ/mol.
@@
The more economical CCSD(T)Λ method yields qualitatively the same answer as CCSDT-3, but not quantitatively.

### v2

@@ 1 Introduction
Theoretical predictions have become an invaluable tool in modern science. As accurate benchmarks in quantum chemistry become more and more common and numerous,1 some questions have arisen about the applicability of the underlying benchmark methods. For increasingly large systems, there are growing discrepancies2 between the interaction energies of large molecules for the main benchmark method of quantum chemistry, which is CCSD(T) (coupled cluster3 including all single and double substitutions and quasiperturbative triple excitations4) and the main benchmark method of solid state materials science, which is QMC (quantum Monte Carlo).5, 6, 7, 8.
@@
The polarizability will increase with the system size of aromatic molecules, and so will the strength of their van der Waals interactions with each other. This polarizability may thus be increasingly hard to calculate with progressively larger basis sets. For example, even small molecules and atoms with huge polarizabilities need many diffuse functions in order to accurately describe these quantities.9
@@
Larger intermolecular interactions are possibly overestimated by the use of perturbative triples in the (T) part, as second-order Møller-Plesset perturbation theory (MP2) strongly overestimates10 intermolecular interactions of aromatic molecules. Various papers have proffered this argument, showing that post-CCSD(T) contributions should become relevant at some point.11, 12, 13 However, one has to bear in mind that while MP2 overestimates interaction energies of stacks, MP3 overcorrects and underestimates (which has inspired the MP2.5 method14). CCSD can be seen as a summation of MP singles, doubles, and disconnected products thereof to infinite order, and underestimates interaction energies as it neglects connected triples (T), which make their first appearance in fourth order. In contrast, however, other post-Hartree-Fock or post-DFT methods such as RPA (random phase approximation15) and DFT-SAPT (density functional theory-based symmetry-adapted perturbation theory16) are argued to yield better, if not correct intermolecular interaction energies of larger dimers.11, 12
@@
For intermolecular interactions with medium to large basis sets, the exact basis set limit is usually bracketed by calculations excluding and including counterpoise corrections,17, 18 with very few notable exceptions, like, for example, the formic acid dimer.19 Thus, an increasingly large gap between these two calculations (including and excluding a basis set correction) would indicate a volatility of a certain post-Hartree-Fock method. Basis set incompleteness errors can also often be estimated using localized orbital or other reduced-scaling approaches.
@@
While the use of CCSD(T) gradients is very often computationally prohibitive, several structures can be considered in order to get a more complete picture of the interaction. This is usually done by just varying the distances, e.g. extending the S66 set20 of intermolecular interactions to the S66x8 set,21 which considers each complex at eight different multiples of the intermonomer separation (at fixed intramonomer geometries), for a total of 528 data points.
@@
Post-Hartree-Fock methods such as coupled cluster theory3 have a clear hierarchy. For example, when static correlation effects are present and not captured well by CCSD(T), one may attempt to walk up the cluster expansion staircase to CCSDT and CCSDT(Q), which will yield some insight about whether ‘the gold standard’ is really that golden.22, 23, 24
@@ 2 Computational Details
The calculations have been performed with various programs, depending on the method employed. The density functional theory (DFT) 25, 26 geometry optimizations utilizing the -B97M-V 27 functional as well as the DLPNO (Domain-Based Local Pair Natural Orbital)-CCSD(T) 28, 29 calculations have been performed with ORCA 30 , DFT-SAPT31, 32 with MOLPRO33, MP334 with QChem35 because of their RI implementation36, and all dRPA15, 37, LNO (Local Natural Orbitals)-CCSD(T)38, LNO-CCSDT(Q)39, and canonical CCSD(T)4 / CCSDT(Q)40 calculations with the MRCC41 program suite. All LNO-CCSD(T) calculations have been using their respective MP2 basis set corrections as implemented in the MRCC program. While some canonical post-CCSD(T) calculations were also carried out using MRCC, the most demanding ones were performed using a development version of the CFOUR program system42.
@@
For the L7 (i.e., seven large noncovalent complexes) set of systems,43 geometries from the original reference have been utilized as-is. All other structures were optimized with the -B97M-V functional27 using the QZVPP44 basis set with full symmetry. For example, in the case of the benzene dimer, the D6h symmetric saddle point has been utilized. This implies that all structures are directly on top of each other, as only their interaction energy matters. They are not parallely displaced, such as the minimum structures of the benzene45 and coronene dimers46 — they are not minima at all. Despite this, the monomer geometries have either been fixed or relaxed, staying within their respective symmetry for the ethylene stacks, whereas for the acene stacks, we have employed completely relaxed geometries. The rationale for this approach is to be able to use as much symmetry as possible, since only the slope of interaction energies of increasingly large dimers matters.
@@
Throughout this study, Dunning’s correlation consistent basis sets with and without diffuse functions, denoted cc-pVZ47 and aug-cc-pVZ48, have been put to use and will sometimes be abbreviated by DZ, TZ, QZ, and 5Z, as well as aTZ, aQZ, and a5Z. The notations {T,Q}Z and {Q,5}Z refer to basis set extrapolations from TZ and QZ or QZ and 5Z basis sets according to the familiar extrapolation formula, respectively.49
@@
Unless specifically indicated otherwise, interaction energies have been counterpoise (CP) corrected by the standard Boys and Bernardi procedure.50
@@
In order to shed additional light on the performance of CCSD(T)4 vs. fully iterative CCSDT51 and CCSDT(Q),40 we carried out additional calculations using two post-CCSD(T) methods that still only have CPU time scaling with system size N: CCSD(T)Λ 52, 53 and CCSDT-254, 51, 55. Cremer and coworkers56 presented a detailed perturbation theoretical analysis of the differences between full CCSDT and various approximations to it. The CCSDT-3 method omits the coupling term from the full CCSDT amplitudes equations, while CCSDT-2 additionally omits all coupling terms from the said equations. Both reduce the CPU time scaling from to at the expense of an energy difference starting at fifth order in perturbation theory:
@@
where it should be noted56 that is the Hermitian conjugate of .
@@ 3.1 L7 Set and the discrepancies between two reference methods
The original starting point of this study was, in fact, the discrepancy between the two reference methods CCSD(T) and FN-DMC (fixed node diffusion Monte Carlo). For this purpose, it is useful to reiterate the differences between them. In addition, we have also performed DFT-SAPT calculations for all these systems. According to the theoretical background of references 11, 12 and point one above, DFT-SAPT may be even more accurate than CCSD(T) for the interaction energies of ever larger systems. In Table 1, we show the DFT-SAPT together with the literature values of various post-Hartree-Fock methods and FN-DMC for four molecules of the L7 set43 as well as the buckycatcher.
@@
| Benzene2 | 10.02 | 7.220 | 6.160 | 11.861 | 11.22 | 13.420 | 19.720, 19.760 |
| GCGC | 52.02 | 36.043 | 35.760 | 54.7 | 57.02 | 56.143 | 74.743, 73.6,62 74.960 |
| C3A | 62.42 | 34.143 | 47.860 | 62.7 | 69.12 | 74.743 | 113.043, 109.2,62 109.760 |
| C2C2PD | 75.62 | 27.743 | 58.260 | 81.5 | 86.12 | 95.443 | 159.943, 154.3,62 155.760 |
| C3GC | 101.32 | 61.843 | 80.860 | 104.6 | 120.02 | 127.243 | 188.843, 182.2,62 183.760 |
| C60CPPA | 130.52 | 93.860 | 157.8 | 174.62 | 354.960 |
@@
It is apparent from these numbers that MP2 often overestimates the interaction energies of large conjugated molecules, whereas MP3 underestimates them. MP2.5,63 which relies on compensation between the errors of MP2 and MP3, still somewhat overestimates the interaction energies when going to larger molecules.64 The values obtained by DFT-SAPT are pretty much exactly between CCSD(T) and fixed-node diffusion Monte-Carlo (FN-DNC) values.2 If we plot interaction energies vs. number of atoms, we obtain values ranging from 0.79 (MP3) to 0.94 (FN-DMC). The slope of FN-DMC is with 1.09 kJ/mol per atom much smaller than the slope of CCSD(T) with 1.46 kJ/mol per atom- with DFT-SAPT and also MP2.5 (1.28 and 1.34 kJ/mol per atom, respectively) in this case being somewhat smaller than CCSD(T).
@@ 3.2 Intermolecular Interaction Energies of Larger Molecules
The first question is: Can we estimate these numbers also from interaction energy slopes of other molecules? In the case that they do, they may offer a hint about the accuracy of different post-Hartree-Fock methods for the interaction energies of large molecules. The systems that may come in mind are the benzene, coronene, and circumcoronene dimers, as the coronene dimer with 72 atoms was already present in the original L7 set. The parallel displaced coronene dimer is yielding a significant difference between FN-DMC with an interaction energy of 75.6 kJ/mol and CCSD(T) with an interaction energy of 86.1 kJ/mol, which is more than 10 kJ/mol. Given, however, that the circumcoronene dimer already has 144 atoms, we need other systems as touchstones for our hypothesis: that we can an interpolate the behavior for really large systems from the slopes of interaction energies of polyaromatic dimers.
To stay as close to the aromatic behavior that we see in the above sequence, we chose the acene dimer series, thus again beginning with the interaction energies of the benzene through hexacene dimers. Starting from heptacene dimer, however, the monomers will undergo cyclodimerization,65, 66, 67 which means that if we want to relax geometries like in the L7 paper (even with symmetry constraints) we will have to stop at hexacene dimer. The latter, with 82 atoms, is about the same size as coronene dimer.
@@ 3.3 Basis Set Effects
First, we try to evaluate the basis set effects of the different methods. If, for example, polarization would be a major issue, the dissociation energies along a chain dimer sequence would not be on a straight line, but rather be flat, as the increasing polarization would be less well described by a small number of basis functions. The number of basis functions, however, also increases linearly with system size, yielding almost perfectly straight lines for different acene series, as shown in Figure 1a-c for LNO-CCSD(T) using tight settings. As counterpoise corrections for LNO and similar methods in small molecules are problematic owing to screening out a substantial part of the ”ghost atoms” contribution [see, e.g., Ref.68], we had to exclude the first couple of cp-corrected interaction energies (for example, the benzene and anthracene dimer for the DZ basis set and the benzene, anthracene and naphthalene dimer for the TZ basis set) in each graph.
@@ 3.4 Local Correlation
Having established the basis set effects, we can estimate canonical CCSD(T) at the basis set limit. Furthermore, especially with the advent of using local correlation methods and extrapolating them,69, 70, 71, 72 we can evaluate their accuracy when going to larger molecules. For this purpose, we compare these methods, especially for a small basis set of double-zeta quality. Again, all of the investigated series exhibit nearly perfect linear regression with an value of 0.996 for methods.
@@ 3.5 Electronic Structure Methods
Here, we investigate several post-Hartree-Fock methods, such as MP2, MP2.5, MP3, the dispersion contribution to DFT-SAPT, as well as dRPA. From these methods, DFT-SAPT and dRPA in particular are deemed to be as more accurate than perturbation theory for large molecules.11, 12 For the acene series, we obtain the trend of MP3 CCSD DFT-SAPT dRPA CCSD(T) MP2.5 MP2. This is hardly surprising, as MP2 very much overestimates the interaction energies of large aromatic molecules, while MP3 underestimated them, as does CCSD. Of course, the question arises about the accuracy of CCSD(T), dRPA, and DFT-SAPT.
For the ethylene stack series, we obtain the trend CCSD MP3 DFT-SAPT dRPA CCSD(T) MP2.5 MP2. As observed in the acene series, MP2 overestimates dispersion contributions while both CCSD and MP3 underestimate them. dRPA yields results similar to CCSD(T), which is clearly observed for the cc-pVZ (X=D,T,Q) basis sets. (We note73 that RPA is equivalent to a ring-diagram simplification of coupled cluster with all doubles, CCD.)
@@ 3.6 Effects of the Geometry
So far, we have used geometries for both acenes and ethylene stacks in which the distance between the two monomers were fully optimized within its (high) symmetry. In order to show that the linear behavior is not an artefact, it may be worth to investigate another geometry with a different intermonomer distance. Here, we chose the stacks at the rather large optimized distance of the ethylene dimer, as shorter distances (e.g. at the hexatetraene distance) would result in a positive interaction energy for the ethylene dimer. The ethylene stack series with relaxed geometry exhibits a slope of 8.70 0.27 kJ/subunit. In contrast, the same series with the fixed geometry at the minimum ethylene dimer distance of 4.46 Å shows a much smaller slope of 1.26 0.24 kJ/subunit, where the uncertainty is about one-fifth of the actual value.
@@ 3.7 Best Estimates of Slopes
Usually, the best estimates in energies are obtained by going to ever larger basis sets and then just extrapolating the energies to the complete basis set (CBS) limit. In our case, however, we can either calculate ever larger systems along the sequences studied with a smaller basis set or increase the basis set size thus adding one dimension to the problem. Our results are summarized in Tables 2 and 3. The slopes obtained from just the first two acene dimers, namely benzene2 and naphthalene2, barely differ compared to those of the whole series up to pentacene; the cp-corrected DZ CCSD(T) slope changes by 0.26 kJ/subunit, and the cp-uncorrected DZ CCSD(T) slope even less at 0.17 kJ/subunit. In general, including larger systems in the fits increases the slopes. From these values, including the cc-pV{T,Q}Z basis set extrapolation which is still feasible for canonical CCSD(T) for the benzene and naphthalene dimer, we can estimate the canonical CCSD(T)/cbs values for the benzene-pentacene series, which will be -23.14-0.26=-23.40 kJ/subunit for the cp-corrected CCSD(T) results and -23.09-0.17=-23.26 kJ/subunit for cp-uncorrected CCSD(T). Finally, this values can then be extended to the hexacene dimer with an aug-cc-pV{Q,5}Z basis set extrapolation.
@@
Of more interest are the post-CCSD(T) slopes. The CCSDT and CCSDT(Q) results for acene dimers (benzene and naphthalene dimers) included in Tables 2 and 3 were obtained using rank-reduced coupled cluster methods described in Refs. 74, 75, 76. In these calculations, the coupled cluster equations are solved within a certain excitation subspace instead of the full space as in the canonical methods. This allows for reducing the costs of the calculations, but introduces an additional parameter (size of the excitation subspace) as a parameter in the rank-reduced formalism. Therefore, it is important that the results are stable with respect to the adopted value of this parameter. In the Supporting Information, we provide a detailed analysis of the convergence of the rank-reduced results with respect to size of the excitation subspace for benzene dimer and naphthalene dimer, showing that the numerical values reported here are sufficiently well converged for the present purposes. The only additional approximation employed in the rank-reduced calculations is the Cholesky decomposition of two-electron integrals. We used the full pivoting variant of this decomposition with a VeryTight threshold () for the diagonal elements. The errors resulting from the decomposition with these settings impact the calculated interaction energies by no more than a few thousands of kJ/mol. The CCSDT method reduces the slopes (in absolute value) by a sizeable 0.75 (cp-corrected DZ) and 0.78 (standard DZ) kJ/subunit. However, CCSDT(Q) (again, in absolute value) reverts this by 0.4 in both instances, shrinking the CCSD(T) to CCSDT(Q) differences in the slopes to 0.35 (cp-corrected DZ) and 0.38 (uncorrected DZ) kJ/subunit.
(We note in passing that the (Q) contribution found here for benzene dimer, 0.50 kJ/mol, is just over half of the value of Karton and Martin77, 0.187–0.200 kcal/mol, or 0.78–0.84 kJ/mol. The latter was obtained from a thermochemical cycle combined with a CCSDT(Q) calculation in an unpolarized double-zeta basis set, and serves as a cautionary tale in this regard.)
@@
For the Hartree-Fock exchange part, this is rather simple, as all calculations were performed at the largest basis sets and the cp-corrected and cp-uncorrected slopes are 11.95 and 11.96 kJ/subunit, respectively, yielding exchange-correlation slopes of -10.5 0.2 kJ/mol per acene ring dimer. For the largest acene dimer hexacene, this would make an interaction energy of approximately 63.0 kJ/mol, considering our best estimates. This value, however, is neglecting the intercepts of 3.3 for Hartree-Fock exchange and 1.9 for correlation, arriving at 57.8 1.4 kJ/mol. CCSD(T), for comparison, would yield slightly larger slopes of -11.1 and -10.9, with an interaction energy of 66.0 for the hexacene dimer. Including the intercepts (3.3 and 2.0), we arrive at a final CCSD(T) energy of 60.7 1.5 kJ/mol.
@@
The more economical CCSD(T)Λ method yields qualitatively the same answer as CCSDT-3 and CCSDT-2, but not quantitatively.